# Distributed Inter-Strand Coupling Current Model for Finite Element Simulations of Rutherford Cables

### v1

@@ Abstract
In this paper, we present the Distributed Inter-Strand Coupling Current (DISCC) model. It is a finite element (FE) model based on a homogenization approach enabling efficient and accurate simulation of the transient magnetic response of superconducting Rutherford cables without explicitly representing individual strands. The DISCC model reproduces the inter-strand coupling current dynamics via a novel mixed FE formulation, and can be combined with the Reduced Order Hysteretic Magnetization (ROHM) and Flux (ROHF) models applied at the strand level in order to reproduce the internal strand dynamics: hysteresis, eddy, and inter-filament coupling currents, as well as ohmic effects. We first analyze the performance of the DISCC model alone, as a linear problem. We then extend the analysis to include the internal strand dynamics that make the problem nonlinear. In all cases, the DISCC model offers a massive reduction of the computational time compared to conventional fully detailed FE models while still accounting for all types of loss, magnetization and inductance contributions. Rutherford cables homogenized with the DISCC model can be directly included in FE models of magnet cross-sections for efficient electro-magneto-thermal simulations of their transient response. We present two possible FE formulations for the implementation of the DISCC model, a first one based on the $h$-$\phi$-formulation, and a second one based on the $h$-$\phi$-$a$-formulation, which is well suited for an efficient treatment of the ferromagnetic regions in magnet cross-sections.
  • October 28, 2025
@@ 1 Introduction
Under transient electric and magnetic excitations, superconducting magnets exhibit loss [1], field distortions [2], and differential inductance [3, 4]. These effects must be simulated accurately to compute the load on cryogenic systems, the temperature and stability margins, and the field errors. Moreover, an accurate description of the transient effects and a detailed understanding of their dynamics is crucial for the design of efficient and robust quench protection systems, in particular for novel quench protection ideas that rely on inducing a high current or field change rate in magnets in order to rapidly transition a large part of them to the normal state [5, 6, 7].
@@
Figure 1: Sequence of models leading to a magnet cross-section with homogenized cables. (b-c-e-f) Illustrative solutions with external field and transport current excitation. Dashed arrows from (a) to (b) and from (d) to (e) represent the reduction of the 3D periodic structure of strands and cables to the 2D models with the CATI method as done in [14] and [15], respectively. Solid orange arrows from (b) to (c) and from (e) to (f) represent homogenization steps, described in [16] and this paper, respectively. Solid black arrows from (c) to (e) and from (f) to (g) illustrate the inclusion of a homogenized model into a larger-scale model.
A number of numerical models describing the IS coupling dynamics in Rutherford cables already exist, they involve lumped elements or continuum models, such as described in [17, 18, 19, 20, 21]. These models usually have low computational costs and can be combined with analytical models describing strand magnetization and IF coupling [22, 23, 24]. However, these approaches require simplifying assumptions, such as neglecting many details about the conductor geometries or the non-uniform external field distribution, neglecting hysteresis effects, and neglecting coupling between different loss mechanisms. For frequencies (or rates of field change) at which different loss contributions are strongly coupled, these methods could potentially result in lower accuracy. Moreover, the above-mentioned models cannot be easily implemented in FE frameworks.
@@
In this paper, we present a novel homogenized 2D FE model for Rutherford cables. This model includes all loss and magnetization contributions, and accounts for their interplay. The original contribution of this paper is a method that reproduces IS coupling current dynamics via a Distributed Inter-Strand Coupling Current (DISCC) model. Besides, the homogenized cable model reproduces the other magnetization and loss mechanisms (hysteresis, eddy, IF, and ohmic) associated with external field and transport current, via distributed versions of the Reduced Order Hysteretic Magnetization (ROHM) and Flux (ROHF) models [16, 27]. In this paper, the effect of an axial field (parallel to the longitudinal cable direction) is not considered, as the model is intended for magnets in which the transverse field dominates. Also, for simplicity, thermal effects are not considered. The temperature is assumed uniform and constant.
The homogenized cable model is built upon a series of recent contributions focusing on the modelling of low-temperature superconducting (LTS) strands and cables. This paper completes the homogenization sequence summarized in Fig. 1. In the homogenized cable model, parameters needed for the DISCC, ROHM, and ROHF models are deduced from reference solutions obtained with detailed strand and cable models using the CATI method [14, 15], see Fig. 1(b) and Fig. 1(e), respectively.
@@
The structure of the paper is summarized in Fig. 2. We start in Section 2 by summarizing the different loss and magnetization mechanisms in superconducting Rutherford cables. Then, in Section 3, we describe a reference cable model with homogenized strands modelled individually, see Fig. 1(e). This model is based on [15], in which the IS coupling currents are modelled with the CATI method [14, 15]. The model is then extended by including strand homogenization with ROHM and ROHF models to account for the other loss and magnetization mechanisms. This model serves as a reference for the homogenized cable model, which we present in Section 4, see Fig. 1(f). The homogenized cable model features DISCC as well as distributed versions of ROHM and ROHF models. We successively describe the strong form of the problem, derive a weak form of it based on the $h$-$\phi$-formulation, and discuss its discretization and implementation in an FE framework. Finally, we discuss the accuracy of the approach with a verification analysis. First, in Section 5, we focus on the DISCC model in a linear setting. Then, in Section 6, we include all cable dynamics in a nonlinear setting and we analyze the results with uniform and non-uniform field excitations, and we verify the calculated differential inductance. Finally, in Section 7, we consider a stack of cables on which we apply the $h$-$\phi$-$a$-formulation, a mixed formulation suited for systems containing ferromagnetic materials, that can be used as an alternative to the $h$-$\phi$-formulation.
@@
The application of the homogenized cable model in a magnet cross-section and its validation with experimental measurements will be the subject of a future paper.
@@
The homogenized cable model, containing the DISCC, ROHM, and ROHF models, can be implemented in any FE tool that allows for user-defined formulations and constitutive laws. In this work, all the models are implemented in open-source and free-to-use software. The FE models are implemented in GetDP [28] within FiQuS (Finite Element Quench Simulator) [29]. FiQuS is developed at CERN as part of the STEAM framework [30]. Geometries and meshes are performed by Gmsh [31].
@@ 2 Loss and magnetization mechanisms
The Rutherford cable can be subject to a time-varying transverse magnetic field and transport current, which induce a non-uniform current density distribution. This current density distribution causes field distortion, differential inductance, and power loss. The power loss is directly associated with the overall current density distribution, but for easier interpretation, it is convenient to split the loss into distinct contributions [1], each associated with a different physical mechanism, see Fig. 3.
@@
The different loss contributions influence each other [1, 14] and their interplay makes the calculation of the total loss with analytical models particularly challenging to perform. In the next section, we introduce a reference Rutherford cable model that allows for a numerical description of the different loss components, accounting for their interplay. It is used as a reference model for the homogenized cable model of Section 4.
@@ 3 Reference cable model
In this section, we present the reference cable model (see Fig. 1(e)) that serves as a reference for the homogenized cable model. We start in Section 3.1 with a linear model reproducing the IS coupling currents, based on the CATI method introduced in [15]. Then, in section 3.2 we complete the model by including the ROHM, ROHF, and current sharing models in order to account for field dynamics inside the strands.
@@ 3.1 Linear model for IS coupling currents
We consider a Rutherford cable made of $N_{\text{s}}$ strands, compressed together and transposed with a transposition length $p_{\text{s}}$. The geometry is periodic along $\hat{\boldsymbol{e}}_{z}$ and has a periodicity length $\ell_{\text{s}}=p_{\text{s}}/N_{\text{s}}$, see Fig. 4. The cable carries a transport current $I_{\text{t}}$ and is subject to a transverse magnetic field (perpendicular to $\hat{\boldsymbol{e}}_{z}$). We assume that the applied magnetic field is independent of $z$, and that there is no axial magnetic field (parallel to $\hat{\boldsymbol{e}}_{z}$).
Figure 4: Reference cable model for $N_{\text{s}}=10$ strands. Strands in (a) are represented without adjacent contacts for better visualization. View (b) is a cross-section of the cable, with FE mesh, along one of the green dash-dotted lines in (a). The electrical circuit in (c) defines connections between axial and transverse currents, with adjacent ($R_{\text{a}}$) and crossing ($R_{\text{c}}$) contact resistances, these connections are assumed to be concentrated in a cross-section of the cable along one of the orange dash-dotted lines in (a). Arrows show the convention for positive currents.
Notations for the cable properties are gathered in Table 1. The table also contains notations for the homogenized model that are introduced later in Section 4.
@@
The CATI method for the cable was introduced in [15], as an extension of the CATI method for strands [14], which itself was inspired by the work in [35]. The CATI method for the cable solves for the axial current density and transverse magnetic field using a 2D FE formulation, e.g., on the mesh of Fig. 4(b). In this 2D FE formulation, the strand currents and voltages are coupled with the IS coupling currents and voltages via circuit equations that encode the periodicity of the cable. The circuit coupling is illustrated in Fig. 4(c). Compared to the circuit presented in [15], the crossing contacts are now grouped into a single connection per strand, hence the lumped resistance value $R_{\text{c}}/2$ for the crossing resistors. This choice has a minor influence on the numerical results and leads to more accurate solutions for excitations in the $\hat{\boldsymbol{e}}_{x}$ direction [^2] With this modification, the transverse currents can be interpreted as flowing midway between two successive - and identical, by periodicity - cross-sections where the axial currents are solved, as is the case for the CATI method applied on the strand [14]..
In [15], strands are modelled as massive conductors [36] with low resistivity. This was necessary to verify the model with a conventional 3D FE reference model, in which a massive conductor approach is required to allow for IS coupling currents.
Here, by contrast with [15], the strands are homogenized. This necessitates modelling them as stranded conductors [36], in which the current density is uniform. In reality, the current density distribution is not uniform due to the composite structure of the strands, eddy currents, IF coupling currents, and hysteresis currents dynamics. However, this non-uniformity is not described explicitly in the model: only the average current density is represented as a uniform field, whereas non-uniformity effects on magnetization, inductance and loss are handled carefully by the ROHM and ROHF models [16], as described in Section 3.2.
We use the $h$-$\phi$-formulation [37, 38] to solve for the cable response and axial currents, because it is well suited to handle the nonlinear functions introduced in Section 3.2 [39, 16]. The numerical domain is denoted by $\Omega$, the strands domain by $\Omega_{\text{s}}$, and a coating domain around the strands by $\Omega_{\text{coat}}$. The coating domain enables straightforward implementation of the CATI method and clear separation of individual strand currents from the total transport current, as was discussed in [15]. Its resistivity $\rho_{\text{coat}}$ is fixed to a sufficiently high value to have a negligible effect on the numerical solution. The conducting domain $\Omega_{\text{c}}$ is equal to $\Omega_{\text{s}}\cup\Omega_{\text{coat}}$, and its complementary non-conducting domain is denoted by $\Omega_{\text{c}}^{\text{C}}$.
@@
The functional $\mathcal{I}_{i}(\boldsymbol{h}^{\prime})$ gives the circulation of $\boldsymbol{h}^{\prime}$ around strand $i\in\{1,\dots,N_{\text{s}}\}$, which is the net current $I_{i}$ (A) flowing in that strand for the (test) function $\boldsymbol{h}^{\prime}$. The associated voltage difference accumulated over a length $\ell_{\text{s}}$ is denoted by $V_{i}$ (V). The functional $\mathcal{I}_{\text{t}}(\boldsymbol{h}^{\prime})$ gives the circulation of $\boldsymbol{h}^{\prime}$ around the whole cable, i.e., the transport current $I_{\text{t}}$ (A) for the (test) function $\boldsymbol{h}^{\prime}$, and $E_{\text{t}}$ denotes the associated voltage per unit length (V/m). Currents are handled with cohomology basis functions [15, 40]. Strand currents and voltages are coupled via the electrical circuit shown in Fig. 4(c). Because the strands are modelled as stranded conductors, no electrical resistivity needs to be defined in $\Omega_{\text{s}}$.
@@
Dynamics in individual strands (hysteresis, eddy, and IF coupling currents) are included in Section 3.2. As a first step, the problem defined by Eq. (3.1) enables the analysis of the IS coupling current dynamics and is used as a reference for the DISCC model that are introduced in Section 4.1. As this model is linear, it can be solved in the frequency domain, using complex phasors for the physical fields under harmonic excitations.
As an illustration, we show in Fig. 5 the strand and IS currents induced in a cable with properties given in Table 1 (with $\rho_{\text{coat}}=10^{-7}$ $\Omega$ m) by a harmonic transverse magnetic field oscillating in two different directions. A magnetic field along $\hat{\boldsymbol{e}}_{y}$ mainly induces crossing IS coupling currents while a magnetic field along $\hat{\boldsymbol{e}}_{x}$ mainly induces adjacent IS coupling currents.
Figure 5: Strand currents and IS coupling currents in the cable for a sinusoidal transverse magnetic field excitation of amplitude of $10$ mT in two directions: (top) along $\hat{\boldsymbol{e}}_{y}$ with frequency $10$ Hz, (bottom) along $\hat{\boldsymbol{e}}_{x}$ with frequency $1000$ Hz. Values are those at the time instant when the field is maximum. The results of the reference cable model (with $N_{\text{s}}=36$) with the CATI method are shown. Two different scales are used in the upper plot for better visibility. The legend is the same for both plots. The strand currents and adjacent IS coupling currents are those of the top layer of strands in the cable, as illustrated in the cable schematics in the bottom right corner of the upper plot (the model describes IS coupling currents as flowing in a middle cross-section). Currents in the bottom layer can be deduced by symmetry, see also Fig. 7.

### v2

@@ Abstract
In this paper, we present the Distributed Inter-Strand Coupling Current (DISCC) model. It is a finite element (FE) model based on a homogenization approach enabling efficient and accurate simulation of the transient magnetic response of superconducting Rutherford cables without explicitly representing individual strands. The DISCC model reproduces the inter-strand coupling current dynamics via a novel mixed FE formulation, and can be combined with the Reduced Order Hysteretic Magnetization (ROHM) and Flux (ROHF) models in order to reproduce the effects of internal strand dynamics: hysteresis, eddy, and inter-filament coupling currents, as well as ohmic effects. The DISCC model offers a massive reduction of the computational time compared to fully detailed FE models and still accounts for all types of loss and magnetization contributions. As a result, Rutherford cables homogenized with the DISCC model can be directly included in FE models of magnet cross-sections for efficient electro-magneto-thermal simulations of their transient response. We present two possible FE formulations for the implementation of the DISCC model, a first one based on the $h$-$\phi$-formulation, and a second one based on the $h$-$\phi$-$a$-formulation, which is well suited for an efficient treatment of the ferromagnetic regions in magnet cross-sections.
@@ 1 Introduction
Under transient electric and magnetic excitations, superconducting magnets exhibit loss [1], field distortions [2], and changing inductance [3, 4]. These effects must be simulated accurately to compute the load on cryogenic systems, the temperature and stability margins, and the field errors [2, 5, 6, 7]. Moreover, an accurate description of the transient effects and a detailed understanding of their dynamics is crucial for the design of efficient and robust quench protection systems, in particular for novel quench protection ideas that rely on inducing a high current or field change rate in magnets via external excitation coils in order to rapidly transition a large part of them to the normal state [8, 9, 10].
@@
Figure 1: Sequence of models leading to a magnet cross-section with homogenized cables. (b-c-e-f) Illustrative solutions with external field and transport current excitation. Dashed arrows from (a) to (b) and from (d) to (e) represent the reduction of the 3D periodic structure of strands and cables to the 2D models with the CATI method as done in [17] and [18], respectively. Solid orange arrows from (b) to (c) and from (e) to (f) represent homogenization steps, described in [19, 20] and this paper, respectively. Solid black arrows from (c) to (e) and from (f) to (g) illustrate the inclusion of a homogenized model into a larger-scale model.
A number of numerical models describing the IS coupling dynamics in Rutherford cables already exist, they involve lumped elements or continuum models, such as described in [21, 22, 23, 24, 25, 26, 27]. These models usually have low computational costs and can be combined with analytical models describing strand magnetization and IF coupling [28, 29, 30]. However, these approaches require simplifying assumptions, such as neglecting many details about the conductor geometries or the non-uniform external field distribution, neglecting hysteresis effects, and neglecting coupling between different loss mechanisms. For frequencies (or rates of field change) at which different loss contributions are strongly coupled, these methods could potentially result in lower accuracy. Moreover, the above-mentioned models cannot be easily implemented in FE frameworks.
@@
In this paper, we present a novel homogenized 2D FE model for Rutherford cables. This model includes all loss and magnetization contributions, and accounts for their interaction. It completes the homogenization sequence summarized in Fig. 1. The original contribution of this paper is the Distributed Inter-Strand Coupling Current (DISCC) model, which reproduces IS coupling current dynamics. Besides, the homogenized cable model reproduces the other magnetization and loss mechanisms (hysteresis, eddy, IF, and ohmic) associated with external field and transport current via distributed versions of the Reduced Order Hysteretic Magnetization (ROHM) and Flux (ROHF) models [19, 20], which are involved as local hysteretic and rate-dependent constitutive laws. Parameters of the DISCC, ROHM, and ROHF models are tuned using detailed strand and cable simulations. Efficient and accurate detailed strand and cable models can be obtained using the CATI method [17, 18], see Fig. 1(b) and Fig. 1(e), respectively, which accounts for filament twist and strand transposition using a pair of 2D models coupled by circuit equations.
@@
The structure of the paper is summarized in Fig. 2. We start in Section 2 by summarizing the different loss and magnetization mechanisms in superconducting Rutherford cables: IS, hysteresis, eddy, IF, and ohmic effects. Then, in Section 3, we describe a reference cable model with homogenized strands modelled individually, see Fig. 1(e), based on [18] in which the IS coupling currents are modelled with the CATI method [17, 18]. This reference model is presented in two steps: first in Section 3.1 (linear) as a simple model containing only the IS dynamics, then in Section 3.2 (nonlinear) with the ROHM and ROHF models to account for the other loss and magnetization mechanisms (hysteresis, eddy, IF, and ohmic) in homogenized strands.
The reference cable model serves as a reference for the homogenized cable model, which we present in Section 4, see Fig. 1(f). The homogenized cable model is also presented in two steps: first in Section 4.1 (linear) with the DISCC model reproducing only the IS dynamics, then in Section 4.2 (nonlinear) with distributed versions of the ROHM and ROHF models in addition to the DISCC model. We define the strong form of the problem, derive a weak form based on the $h$-$\phi$-formulation, and discuss its discretization and implementation in an FE framework.
Next, we discuss the accuracy of the approach with a verification analysis. First, in Section 5, we focus solely on the IS dynamics based on the linear models which allow to focus on the DISCC model contribution only. Then, in Section 6, we generalize the verification by considering all the cable dynamics with the nonlinear models. Finally, in Section 7, we consider a stack of cables on which we apply the $h$-$\phi$-$a$-formulation, a mixed formulation suited for systems containing ferromagnetic materials, that can be used as an alternative to the $h$-$\phi$-formulation.
@@
In this paper, the effect of an axial field (parallel to the longitudinal cable direction) is not considered, because the model is intended for magnets in which the transverse field dominates. Also, for simplicity, thermal effects are not considered and temperature is assumed uniform and constant. The application of the homogenized cable model in a magnet cross-section and its validation with experimental measurements will be the subject of a future paper.
@@
The homogenized cable model, containing the DISCC, ROHM, and ROHF models, can be implemented in any FE tool that allows for user-defined formulations and constitutive laws. In this work, all the models are implemented in open-source and free-to-use software. The FE models are implemented in GetDP [33] within FiQuS (Finite Element Quench Simulator) [34]. FiQuS is developed at CERN as part of the STEAM framework [35]. Geometries and meshes are performed by Gmsh [36]. The input files for the simulations are provided online [^2] https://gitlab.cern.ch/steam/analyses/discc-model. Results were obtained with CERNGetDP version 2025.12.1 and FiQuS version 2025.12.0.
@@ 2 Loss and magnetization mechanisms
The Rutherford cable can be subject to a time-varying transverse magnetic field and transport current, which induce a non-uniform current density distribution. This current density distribution causes field distortion [2], changing inductance [3, 4], and power loss [1]. The total power loss $p_{\text{tot}}$ is directly associated with the overall current density distribution, but for easier interpretation, it is convenient to split the loss into distinct contributions [1], each associated with a distinct physical mechanism, see Fig. 3.
@@
The different loss contributions influence each other [1, 17] and their interaction makes the calculation of the total loss with analytical models particularly challenging to perform. In the next section, we introduce a reference Rutherford cable model that allows for a numerical description of the different loss components, accounting for their interaction. It is used as a reference model for the homogenized cable model of Section 4.
@@ 3 Reference cable model
In this section, we present the reference cable model (see Fig. 1(e)) that serves as a reference for the homogenized cable model. We start in Section 3.1 with a (linear) model reproducing only the IS coupling current dynamics, based on the CATI method introduced in [18]. Then, in section 3.2 we generalize it by including the ROHM and ROHF models, as well as a current sharing law in order to account for all the field dynamics inside the strands (hysteresis, eddy, IF, ohmic effects).
@@ 3.1 Linear model for IS coupling currents
We consider a Rutherford cable made of $N_{\text{s}}$ strands, compressed together and transposed with a transposition length $p_{\text{s}}$. The geometry is periodic along $\hat{\boldsymbol{e}}_{z}$ and has a periodicity length $\ell_{\text{s}}=p_{\text{s}}/N_{\text{s}}$, see Fig. 4. The cable carries a transport current $I_{\text{t}}$ and is subject to a transverse magnetic field (perpendicular to $\hat{\boldsymbol{e}}_{z}$). We assume that the applied magnetic field is independent of $z$, and that there is no axial magnetic field (parallel to $\hat{\boldsymbol{e}}_{z}$). Notations for the cable properties are gathered in Table 1.
Figure 4: Reference cable model for $N_{\text{s}}=10$ strands. Illustration of the connections between strands. Strands in (a) are represented without adjacent contacts for better visualization. View (b) is a cross-section of the cable, with FE mesh, along one of the green dash-dotted lines in (a). The electrical circuit in (c) defines connections between axial and transverse currents, with adjacent ($R_{\text{a}}$) and crossing ($R_{\text{c}}$) contact resistances, these connections are assumed to be concentrated in a cross-section of the cable along one of the orange dash-dotted lines in (a). Arrows show the convention for positive currents.
@@
The cable is solved with the CATI method described in [18], based on the work in [17, 40]. The CATI method for the cable solves for the axial current density and transverse magnetic field using a 2D FE formulation, e.g., on the mesh of Fig. 4(b), and coupled the strand currents and voltages with the IS coupling currents and voltages via circuit equations that encode the periodicity of the cable. The circuit coupling is illustrated in Fig. 4(c). Compared to the circuit presented in [18], the crossing contacts are now grouped into a single connection per strand, hence the lumped resistance value $R_{\text{c}}/2$ for the crossing resistors. This choice has a minor influence on the numerical results but leads to more accurate solutions for excitations in the $\hat{\boldsymbol{e}}_{x}$ direction [^3] With this modification, the transverse currents can be interpreted as flowing midway between two successive - and identical, by periodicity - cross-sections where the axial currents are solved, as is the case for the CATI method applied on the strand [17]..
In [18], strands are modelled as massive conductors [41] with low resistivity. This was necessary to verify the model with a conventional 3D FE reference model, in which a massive conductor approach is required to allow for the flow of IS coupling currents.
Here, by contrast with [18], the strands are homogenized and modelled as stranded conductors [41], in which the current density is uniform. In reality, the current density distribution is not uniform due to the composite structure of the strands, eddy currents, IF coupling currents, and hysteresis currents dynamics. However, this non-uniformity is not described explicitly in the model: only the average current density is represented as a uniform, homogenized field, whereas the impact of non-uniformity on magnetization, inductance and loss is handled carefully by the ROHM and ROHF models [19], as described in Section 3.2. Also, the strand resistivity is assumed to be exactly zero for the moment. A nonlinear resistivity describing the current sharing law will be introduced in Section 3.2.
We use the $h$-$\phi$-formulation [42, 43] to solve for the cable response and axial currents. The numerical domain is denoted by $\Omega$, the strands domain by $\Omega_{\text{s}}$, and a coating domain around the strands by $\Omega_{\text{coat}}$. The coating domain enables straightforward implementation of the CATI method and clear separation of individual strand currents from the total transport current, as was discussed in [18]. Its resistivity $\rho_{\text{coat}}$ is fixed to a sufficiently high value to have a negligible effect on the numerical solution. The conducting domain $\Omega_{\text{c}}$ is equal to $\Omega_{\text{s}}\cup\Omega_{\text{coat}}$, and its complementary non-conducting domain is denoted by $\Omega_{\text{c}}^{\text{C}}$.
@@
The functional $\mathcal{I}_{i}(\boldsymbol{h}^{\prime})$ gives the circulation of $\boldsymbol{h}^{\prime}$ around strand $i\in\{1,\dots,N_{\text{s}}\}$, which is the net current $I_{i}$ (A) flowing in that strand for the (test) function $\boldsymbol{h}^{\prime}$. The associated voltage difference accumulated over a length $\ell_{\text{s}}$ is denoted by $V_{i}$ (V). The functional $\mathcal{I}_{\text{t}}(\boldsymbol{h}^{\prime})$ gives the circulation of $\boldsymbol{h}^{\prime}$ around the whole cable, i.e., the transport current $I_{\text{t}}$ (A) for the (test) function $\boldsymbol{h}^{\prime}$, and $E_{\text{t}}$ denotes the associated voltage per unit length (V/m). Currents are handled with cohomology basis functions [18, 44]. Strand currents and voltages are coupled via the electrical circuit shown in Fig. 4(c). Because the strands are modelled as stranded conductors, no electrical resistivity needs to be defined in $\Omega_{\text{s}}$, such that axial currents in $\Omega_{\text{s}}$ do not generate resistive voltage drops.
@@
The problem defined by Eq. (3.1) enables to focus on the IS coupling current dynamics. As this model is linear, it can be solved in the frequency domain, using complex phasors for the physical fields under harmonic excitations. As an illustration, we show in Fig. 5 the strand and IS currents induced in a cable with properties given in Table 1 (with $\rho_{\text{coat}}=10^{-7}$ $\Omega$ m) by a harmonic transverse magnetic field oscillating in two different directions. A magnetic field along $\hat{\boldsymbol{e}}_{y}$ mainly induces crossing IS coupling currents, while a magnetic field along $\hat{\boldsymbol{e}}_{x}$ mainly induces adjacent IS coupling currents.
Figure 5: Strand currents and IS coupling currents in the cable for a sinusoidal transverse magnetic field excitation of amplitude of $10$ mT in two directions: (top) along $\hat{\boldsymbol{e}}_{y}$ with frequency $10$ Hz, (bottom) along $\hat{\boldsymbol{e}}_{x}$ with frequency $1000$ Hz. Values are those at the time instant when the field is maximum. The results of the reference cable model (with $N_{\text{s}}=36$) with the CATI method are shown. Two different scales are used in the upper plot for better visibility. The legend is the same for both plots. The strand currents and adjacent IS coupling currents are those of the top layer of strands in the cable, as illustrated in the cable schematics in the upper plot (the model describes IS coupling currents as flowing in a middle cross-section). Currents in the bottom layer can be deduced by symmetry, see also Fig. 7.